A binomial probability experiment is conducted with the given parameters. Compute the probability of x successes in the n independent trials of the experiment. nequals10​, pequals0.5​, xequals3 Upper P left parenthesis 3 right parenthesisequals nothing ​(Do not round until the final answer. Then round to four decimal places as​ needed.)
Answer #1

Please note nCx = n! / [(n-x)!*x!]

Binomial Probability = nCx * (p)x * (q)n-x, where n = number of trials and x is the number of successes.

Here n = 10, p = 0.05, q = 1 – p = 0.5.

P(X = 3) = 10C3 * (0.5)3 * (0.5)10-3 = 0.1171